# Critical Masses and Numerical Computation of Massive Scalar Quasinormal Modes in Schwarzschild Black Holes

### v1

@@ VI.2 Quasinormal modes
Table 3 presents the quasinormal modes for a massive scalar field with $m=0.1$ as a function of the angular momentum number $l$ and overtone number $n$. Note that, for each fixed $l$, the real part of the fundamental frequency ($n=0$) increases monotonically with $l$, reflecting the higher oscillation frequencies associated with modes of greater angular momentum. This is accompanied by a gradual decrease in the magnitude of the imaginary part, which indicates that higher $l$ modes are longer-lived. This result is consistent with previous studies on massive scalar perturbations [46, 22, 47].
@@
The presence of the scalar field mass manifests as an overall shift in the quasinormal modes spectrum: the real part of the frequencies is slightly increased compared to the massless case, while the imaginary part is reduced in magnitude, leading to longer-lived perturbations. This tendency, observed across all $l$ and $n$, is in agreement with theoretical expectations and previous numerical results [48, 49]. The slower decay of massive field perturbations is particularly relevant for gravitational wave phenomenology, as it implies that massive scalar fields can produce late-time signals with potentially observable imprints [22].
We now analyze the behavior of quasinormal modes (QNMs) as the scalar field mass approaches the critical values $m_{\rm lim}$ and $m_{\rm max}$. We begin with the case $l=0$, for which the critical masses are $m_{\rm lim}=0.385$ and $m_{\rm max}=0.5$. The evolution of the fundamental mode with increasing mass is illustrated in Fig. 3. The real part of the QNM frequency grows with the mass, while the imaginary part decreases and approaches zero at a certain mass value. A notable feature is the qualitative change in behavior observed at $m=0.5$: the real part exhibits a drop followed by a less smooth rise, while the imaginary part initially increases, then decreases, vanishing close to $m\simeq 0.7$. This change occurs when the effective potential ceases to have a distinct peak, which, according to the interpretation of QNMs described in Ref. [30], would indicate the non-existence of QNMs for masses beyond $m_{\rm max}$. Nevertheless, it is possible to obtain numerical frequency values beyond this mass, including cases where the imaginary part vanishes. Modes with vanishing imaginary parts are identified as long-lived modes, corresponding to perturbations with arbitrarily long lifetimes, and have been extensively studied in the literature [50, 22, 51, 52, 43].
@@
For $l=1$, the critical values are $m_{\rm lim}=0.794$ and $m_{\rm max}=0.931$. As shown in Fig. 4, the real part of the QNM frequency again increases with the mass, while the imaginary part decreases as the mass approaches $m_{\rm max}$. A change in behavior is observed at $m=m_{\rm max}$, although it is less pronounced than in the $l=0$ case. Beyond this point, the imaginary part continues to decrease and eventually approaches zero near $m\simeq 1.0$, indicating the emergence of a long-lived mode.
@@
A similar trend is observed for $l=2$, with $m_{\rm lim}=1.276$ and $m_{\rm max}=1.480$. As shown in Fig. 5, the transition in the real and imaginary parts near $m_{\rm max}$ is even milder, indicating a systematic weakening of this effect with increasing multipole number. Nevertheless, long-lived modes still appear beyond $m_{\rm max}$, with the imaginary part vanishing near $m\simeq 1.6$. This indicates that the emergence of long-lived modes occurs at different values of $l$, although the precise mass at which they appear shifts to higher values as $l$ increases.
@@
In summary, for all considered multipoles, the real and imaginary parts of the QNM frequencies exhibit a clear qualitative transition near $m_{\rm max}$, which corresponds to the disappearance of the peak in the effective potential. This transition may indicate the nonexistence of physical QNMs for $m>m_{\rm max}$. Despite this, both numerical methods continue to yield frequency values beyond $m_{\rm max}$, consistently indicating the emergence of long-lived modes at a mass $m_{zd}$, referred to as the zero-damping mass. At this point, the imaginary part of the frequency vanishes, and the mode becomes effectively non-decaying. While the exact value of $m_{zd}$ differs slightly between the two methods, its existence is robust. By contrast, $m_{\rm lim}$ has no apparent effect on the qualitative behavior of the QNM spectrum. We therefore identify three critical mass scales: $m_{\rm max}$, which governs the qualitative transition in both the real and imaginary parts of the frequencies; $m_{\rm lim}$, which appears to be irrelevant for the spectral dynamics within the explored parameter space; and $m_{zd}$, which is not associated with the effective potential but rather signals the onset of long-lived modes.
Finally, regarding the comparison of the two methods in the massive case, we observe that both approaches yield consistent results up to mass values close to $m_{\rm max}$. Beyond this point, small differences appear in the real and imaginary parts of the frequencies, the most notable being the precise value of $m_{zd}$. Importantly, this discrepancy diminishes as the multipole number $l$ increases. We also verified the numerical stability of both methods with respect to the matrix size $N$ and the number of terms $j$, considering values $N=j=100$, 500500, and $1000$. In all cases, the results remain unchanged, demonstrating the high stability and reliability of both approaches.
@@ VII Conclusions
In this work, we have presented a comprehensive analysis of the quasinormal modes (QNMs) of a massive scalar field in Schwarzschild spacetime, employing two complementary numerical techniques: the Hill-determinant method and Leaver’s continued-fraction method. Both approaches proved robust and yielded consistent spectra across a wide range of field masses and angular momenta.
A central result of our study is the identification of three critical mass thresholds, $_{\rm lim}, mm$$_{\rm max}$, and $m_{zd}$, which govern qualitative changes in the QNM spectrum. As the scalar field mass approaches $m_{\rm max}$, the effective potential loses its characteristic peak, and the frequencies undergo a marked transition in behavior. Beyond this threshold, following the classical interpretation of Ref. [30], physical QNMs no longer exist. Nevertheless, our numerical analysis showed that there are solutions that evolve into long-lived, or zero-damping, states at a distinct value of the mass, denoted $m_{zd}$, where the imaginary part of the frequency vanishes and the mode becomes essentially non-decaying. Although the precise value of $m_{zd}$ shows minor discrepancies between the two methods, its presence is robust across all multipoles and shifts to larger values with increasing $l$.
@@
In summary, our study provides a detailed numerical characterization of massive scalar QNMs, clarifies the role of mass thresholds in shaping their spectra, and demonstrates the complementary strengths of the Hill-determinant and continued-fraction methods. Future extensions of this framework could include the analysis of rotating or charged black holes, as well as possible connections with quasi-bound states.
Acknowledgements
MFSA thanks Fundação de Amparo à Pesquisa e Inovação do Espírito Santo (FAPES, Brazil) for support. LGM thanks Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil) for partial financial support—Grant: 307901/2022-0 (LGM). BPP thanks Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) for financial support

### v2

@@ VI.2 Quasinormal modes
Table 3 presents the quasinormal modes for a massive scalar field with $m=0.1$ as a function of the angular momentum number $l$ and overtone number $n$. Note that, for each fixed $l$, the real part of the fundamental frequency ($n=0$) increases monotonically with $l$, reflecting the higher oscillation frequencies associated with modes of greater angular momentum. This is accompanied by a gradual decrease in the magnitude of the imaginary part, which indicates that higher $l$ modes are longer-lived. This result is consistent with previous studies on massive scalar perturbations [Ma:2006by, Zinhailo:2024jzt, Konoplya:2005et].
The quasinormal spectrum of a massive field exhibits distinctive and nontrivial features, first reported in Ref. [1985ApJ...291L..33S]. In contrast to the comparatively simple spectra of massless fields, the presence of mass introduces characteristic structures in the frequency domain. As the field mass parameter $m$ increases, the damping rate of the fundamental (least damped) mode decreases steadily, approaching zero at a critical value of $m$. Beyond this threshold, the fundamental mode disappears from the spectrum, and the first overtone assumes dominance. This pattern repeats for progressively larger $m$, with successive overtones undergoing the same transition. Consequently, the spectrum of massive fields may contain quasinormal modes with arbitrarily small damping, corresponding to extremely long-lived oscillations that closely resemble standing waves.
@@
The presence of the scalar field mass manifests as an overall shift in the quasinormal modes spectrum: the real part of the frequencies is slightly increased compared to the massless case, while the imaginary part is reduced in magnitude, leading to longer-lived perturbations. This tendency, observed across all $l$ and $n$, is in agreement with theoretical expectations and previous numerical results [Becar:2023zbl, Lu:2023par]. The slower decay of massive field perturbations is particularly relevant for gravitational wave phenomenology, as it implies that massive scalar fields can produce late-time signals with potentially observable imprints [Zinhailo:2024jzt].
We now analyze the behavior of QNMs as the scalar field mass approaches the critical values $m_{\rm lim}$ and $m_{\rm max}$. We begin with the case $l=0$, for which the critical masses are $m_{\rm lim}=0.385$ and $m_{\rm max}=0.5$. The evolution of the fundamental mode with increasing mass is illustrated in Figure 3. The real part of the QNM frequency grows with the mass, while the imaginary part decreases and approaches zero at a certain mass value. A notable feature is the qualitative change in behavior observed at $m=0.5$: the real part exhibits a drop followed by a less smooth rise, while the imaginary part initially increases, then decreases, vanishing close to $m\simeq 0.7$. This change occurs when the effective potential ceases to have a distinct peak, which, according to the interpretation of QNMs described in Ref. [1985ApJ...291L..33S], would indicate the nonexistence of QNMs for masses beyond $m_{\rm max}$. Nevertheless, it is possible to obtain numerical frequency values beyond this mass, including cases where the imaginary part vanishes. Modes with vanishing imaginary parts are identified as long-lived modes, corresponding to perturbations with arbitrarily long lifetimes, and have been extensively studied in the literature [Ohashi:2004wr, Zinhailo:2024jzt, Konoplya:2017tvu, Konoplya:2005hr, Konoplya:2004wg].
@@
For $l=1$, the critical values are $m_{\rm lim}=0.794$ and $m_{\rm max}=0.931$. As shown in Figure 4, the real part of the QNM frequency again increases with the mass, while the imaginary part decreases as the mass approaches $m_{\rm max}$. A change in behavior seems to occur at $m=m_{\rm max}$, although the discrepancy between the Leaver and Hill methods makes it unclear whether this feature is real. Beyond this point, the imaginary part continues to decrease and eventually approaches zero near $m\simeq 1.0$, indicating the emergence of a long-lived mode.
@@
As shown in Figure 5, a similar trend is observed for $l=2$, with $m_{\rm lim}=1.276$ and $m_{\rm max}=1.480$. Nevertheless, long-lived modes still appear beyond $m_{\rm max}$, with the imaginary part vanishing near $m\simeq 1.6$. This indicates that the emergence of long-lived modes occurs at different values of $l$, although the precise mass at which they appear shifts to higher values as $l$ increases.
@@
In summary, for the $l=0$ mode, the real and imaginary parts of the QNM frequencies exhibit a clear qualitative transition near $m_{\rm max}$, which corresponds to the disappearance of the peak in the effective potential. A similar, though less evident, behavior is suggested for the $l=1$ and $l=2$ modes. This transition may indicate the nonexistence of physical QNMs for $m>m_{\rm max}$. Despite this, both numerical methods continue to yield frequency values beyond $m_{\rm max}$, consistently indicating the emergence of long-lived modes at a mass $m_{zd}$, referred to as the zero-damping mass. At this point, the imaginary part of the frequency vanishes, and the mode becomes effectively nondecaying. While the exact value of $m_{zd}$ differs slightly between the two methods, its existence is robust. By contrast, $m_{\rm lim}$ has no apparent effect on the qualitative behavior of the QNM spectrum. We therefore identify three critical mass scales: $m_{\rm max}$, which governs the qualitative transition in both the real and imaginary parts of the frequencies; $m_{\rm lim}$, which appears to be irrelevant for the spectral dynamics within the explored parameter space; and $m_{zd}$, which is not associated with the effective potential but rather signals the onset of long-lived modes.
Finally, regarding the comparison between the two numerical schemes in the massive case, we find that both methods yield consistent results up to field masses close to $m_{\text{max}}$. Beyond this value, discrepancies begin to appear in both the real and imaginary parts of the quasinormal frequencies, the most significant being the precise determination of the critical mass $m_{zd}$. Notably, this difference tends to diminish as the multipole number $l$ increases. We have also verified the numerical stability of both approaches with respect to the matrix size $N$ and the number of terms $j$, considering values $N=j=100$, 500500, and $1000$. In all cases, the results remain unchanged, confirming the high stability and reliability of both methods.
This slight divergence between the two approaches originates from their distinct mathematical formulations. In Leaver’s continued-fraction method, the quasinormal frequencies are obtained from the convergence condition of an infinite series whose behavior depends sensitively on the asymptotic structure of the recurrence coefficients. In contrast, the Hill-determinant method relies on truncating a finite matrix and imposing the vanishing of its determinant, thereby avoiding explicit convergence issues inherent to infinite series. A well-known case where the continued-fraction approach may fail is that of purely imaginary quasinormal frequencies, for which the recurrence relation becomes numerically unstable and convergence deteriorates [Cook:2016ngj, Fortuna:2020obg]. In this regard, the Hill-determinant technique generally exhibits fewer convergence pathologies and often remains stable even in such extreme regimes. Nevertheless, this observation alone is not sufficient to determine which of the two methods performs better for the present problem. Despite this, qualitatively, both methods indicate the existence of $m_{zd}$, demonstrating that its presence is a robust feature of the spectrum.
@@ VII Conclusions
In this work, we have presented a comprehensive analysis of the QNMs of a massive scalar field in Schwarzschild spacetime, employing two complementary numerical techniques: the Hill-determinant method and Leaver’s continued-fraction method. Both approaches proved robust and yielded consistent spectra across a wide range of field masses and angular momenta.
A central result of our study is the identification of three critical mass thresholds, $_{\rm lim}, mm$$_{\rm max}$, and $m_{zd}$, which govern qualitative changes in the QNM spectrum. As the scalar field mass approaches $m_{\rm max}$, the effective potential loses its characteristic peak, and the frequencies undergo a marked transition in behavior. Beyond this threshold, following the classical interpretation of Ref. [1985ApJ...291L..33S], physical QNMs no longer exist. Nevertheless, our numerical analysis showed that there are solutions that evolve into long-lived, or zero-damping, states at a distinct value of the mass, denoted $m_{zd}$, where the imaginary part of the frequency vanishes and the mode becomes essentially nondecaying. Although the precise value of $m_{zd}$ shows minor discrepancies between the two methods, its presence is robust across all multipoles and shifts to larger values with increasing $l$.
@@
In summary, our study provides a detailed numerical characterization of massive scalar QNMs, clarifies the role of mass thresholds in shaping their spectra, and demonstrates the complementary strengths of the Hill-determinant and continued-fraction methods. Future extensions of this framework could include the analysis of rotating or charged black holes, as well as possible connections with quasibound states.
ACKNOWLEDGMENTS
M.F.S.A. thanks Fundação de Amparo à Pesquisa e Inovação do Espírito Santo (FAPES, Brazil) for support. L.G.M. thanks Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil) for partial financial support—Grant: 307901/2022-0. B.P.P. thanks Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) for financial support
DATA AVAILABILITY
The data that support the findings of this article are openly available [alves_2025_17580446].